Math Mode Reference Card

Variables

All math mode commands are under the prefix key specified by LaTeX-math-abbrev-prefix, default is ”‘”.

You can define your own math mode commands by setting the variable LaTeX-math-list before loading LaTeX-math-mode.

Greek Letters

α  (alpha)  a
β  (beta)  b
δ  (delta)  d
ϵ  (epsilon)  e
ϕ  (phi)  f
γ  (gamma)  g
η  (eta)  h
κ  (kappa)  k
λ  (lambda)  l
μ  (mu)  m
∇  (nabla)  N
ν  (nu)  n
ω  (omega)  o
π  (pi)  p
θ  (theta)  q
ρ  (rho)  r
σ  (sigma)  s
τ  (tau)  t
υ  (upsilon)  u
χ  (chi)  x
ψ  (psi)  y
ζ  (zeta)  z
Δ  (Delta)  D
Γ  (Gamma)  G
Θ  (Theta)  Q
Λ  (Lambda)  L
Ψ  (Psi)  Y
Π  (Pi)  P
Σ  (Sigma)  S
Υ  (Upsilon)  U
Φ  (Phi)  V
Ω  (Omega)  O

Symbols

→ (rightarrow) C-f ← (leftarrow) C-b

↑  (uparrow)  C-p
↓  (downarrow)  C-n
≤  (leq)  <
≥  (geq)  >

˜ (tilde)

∞ (infty) I

∀  (forall)  A
∃  (exists)  E

(not) !

∈  (in)  i
×  (times)  *
·  (cdot)  .
⊂  (subset)  {
⊃  (supset)  }
⊆  (subseteq)  [
⊇  (supseteq)  ]
\  (backslash)  \
\  (setminus)  /
∪  (cup)  +
∩  (cap)  -
⟨  (langle)  (
⟩  (rangle)  )

exp (exp) C-e sin (sin) C-s cos (cos) C-c sup (sup) C-^ inf (inf) C-_ det (det) C-d lim (lim) C-l tan (tan) C-t

ˆ  (hat)  ^
∨  (vee)  v

Miscellaneous

cal letters c LETTER

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1987 Free Software Foundation, Inc.
Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992 Kresten Krab Thorup
Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 Per Abrahamsen

Permission is granted to make and distribute copies of this card provided the copyright notice and this permission notice are preserved on all copies.